### v4

# Learnability Window in Gated Recurrent Neural Networks

Abstract
We develop a theoretical framework that explains how gating mechanisms determine the learnability window $\mathcal{H}_{N}$ of recurrent neural networks, defined as the largest temporal horizon over which gradient information remains statistically recoverable. While classical analyses emphasize numerical stability of Jacobian products, we show that stability alone is insufficient: learnability is governed instead by the effective learning rates $\mu_{t,\ell}$, per-lag and per-neuron quantities obtained from first-order expansions of gate-induced Jacobian products in Backpropagation Through Time. These effective learning rates act as multiplicative filters that control both the magnitude and anisotropy of gradient transport. Under heavy-tailed ($\alpha$-stable) gradient noise, we prove that the minimal sample size required to detect a dependency at lag $\ell$ scales as $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, where $f(\ell)=\|\mu_{t,\ell}\|_{1}$ is the effective learning rate envelope and $\kappa_{\alpha}=\alpha/(\alpha-1)$ is the concentration exponent governing empirical averages. This yields an explicit characterization of $\mathcal{H}_{N}$ and closed-form scaling laws for logarithmic, polynomial, and exponential decay of $f(\ell)$. The theory shows that the time-scale spectra induced by the effective learning rates are the dominant determinants of learnability: broader or more heterogeneous spectra slow the decay of $f(\ell)$, enlarging the learnability window, while heavy-tailed noise uniformly compresses $\mathcal{H}_{N}$ by slowing statistical concentration to $N^{-1/\kappa_{\alpha}}$. By integrating gate-induced time-scale geometry with gradient noise and sample complexity, the framework identifies effective learning rates as the primary objects that determine whether, when, and over what horizons recurrent networks can learn long-range temporal dependencies.
1 Introduction
Recurrent neural networks (RNNs) are fundamental models for processing sequential data, yet their ability to learn long-range dependencies remains difficult to characterize. While gated architectures such as the LSTM and GRU have greatly improved stability and performance, it is still unclear how gating mechanisms determine which temporal dependencies are actually learnable. Most existing analyses focus on dynamical stability, spectral properties, or mean-field approximations, but provide little insight into the statistical conditions under which information from distant past states remains recoverable during training. A common implicit assumption is that long-range learning is governed primarily by the numerical stability of Jacobian products. However, stability alone is not sufficient: even perfectly stable gradients may be too attenuated or too noisy to contain a statistically usable signal. Thus, traditional dynamical criteria do not answer the central question of when gradients transported by Backpropagation Through Time (BPTT) retain enough information to influence parameter updates.
This work builds on our previous analysis of time-scale coupling between state and parameter dynamics in RNNs [30], where gating mechanisms were shown to induce heterogeneous time scales that shape both state evolution and gradient flow. Here we extend that perspective by developing a quantitative theory of finite-horizon learnability that explicitly connects the structure of gating to the statistics of gradient transport. The central concept is the effective learning rate, denoted $\mu_{t,\ell}$, which measures how BPTT re-weights gradient signals across time for each neuron and lag $\ell$. By expanding gate-induced Jacobians to first order, we obtain closed-form expressions for $\mu_{t,\ell}$ in various gated RNNs, including the LSTM and GRU. These rates show how gating acts as a multiplicative filter that modulates both the magnitude and the directional structure of gradient transport, thereby linking the temporal dynamics of the hidden state to the rate and geometry of parameter updates.
Using this formulation, we develop a statistical theory of finite-horizon learnability under heavy-tailed ($\alpha$-stable) gradient noise. Because learning relies on empirical averages over independent training sequences whose fluctuations may be heavy-tailed, statistical concentration plays a central role. We show that the effective learning rates determine the detectability of lagged dependencies: the minimal number of training samples required to learn a dependency at lag $\ell$ scales as $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, where $f(\ell)=\|\mu_{t,\ell}\|_{1}$ is the effective learning rate envelope and $\kappa_{\alpha}=\alpha/(\alpha-1)$ is the concentration exponent associated with $\alpha$-stable averages. From this relation, we derive an explicit expression for the learnability window $\mathcal{H}_{N}$, defined as the largest temporal horizon over which gradient information remains statistically recoverable. The resulting scaling laws for $\mathcal{H}_{N}$ are determined directly by the decay of $f(\ell)$, quantifying how gate-induced time-scale structure and noise statistics jointly constrain temporal learning.
The theory predicts that broad and heterogeneous time-scale spectra expand the learnability window, while heavy-tailed gradient noise uniformly compresses it by slowing statistical concentration. These predictions are validated by empirical studies on several gated RNN architectures, confirming that the distribution and decay of the effective learning rates capture the essential mechanisms that govern when, and for how long, recurrent networks can exploit past information during training.
2 Related work
The introduction of gating mechanisms in LSTM [18] and GRU [8] was pivotal for controlling temporal credit assignment in recurrent neural networks. Early theoretical works linked the forget gate to an interpretable exponential decay and controllable memory retention [13, 44], while more recent studies connect gating to continuous-time or implicit ODE perspectives [34, 4, 14]. These contributions illuminate how gates shape state dynamics, but do not analyze how gating multiplicative structures affect parameter updates and the capacity for learning long-range dependencies. Our framework bridges that gap by deriving per-lag effective learning rates from products of Jacobians and linking them to finite-horizon learnability.
A complementary line of work models gates as learned rescaling or time-warping mechanisms. Tallec and Ollivier [44] formalized gating as adaptive time dilation, and recent geometry-based analyses [25] study how gating choices influence information flow. We build on these ideas by explicitly casting gate-derived Jacobian terms into effective learning rates and embedding them in detectability bounds via mutual information / Fano theory.
The classical vanishing/exploding gradients problem motivated many stabilization techniques: clipping, spectral regularization, and orthogonality or unitary RNNs [3, 37, 1, 46, 21]. Hierarchical, dilated, or skip architectures shorten propagation paths [23, 5, 9], while continuous-time and neural ODE approaches enhance stability and expressivity [39, 16]. Works on orthogonal / dynamical isometry control the conditioning of gradient transport [6, 38] but do not address whether transported gradients still carry statistically reliable information over long lags. In contrast, we analyze statistical detectability, showing that even when Jacobians remain stable, learning can fail if the effective learning rates decay too quickly.
In the broader learning systems literature, trainability has been studied via mean-field theory, spectral initialization, curvature, and geometry of parameter landscapes [41, 38, 27, 10, 32, 28, 42, 48, 22, 49]. In recurrent and transformer settings, recent work links trainability to Jacobian spectra, Fisher information, and anisotropic learning rates [47, 2, 36, 12]. Our framework is complementary: it provides a finite-horizon, sample-complexity view anchored in gradient statistics, showing how the tail behavior of gradient noise and gate structure jointly determine which lags are learnable.
Simsekli et al. [43] show that mini-batch stochastic gradient descent (SGD) noise often exhibits $\alpha$-stable, heavy-tailed behavior rather than Gaussian tails. Subsequent research has further examined the algorithmic and theoretical consequences of such heavy-tailed gradient statistics. Hübler et al. [19] establish a unified framework connecting gradient clipping and normalization, proving that both act as implicit variance control mechanisms for $\alpha$-stable noise and can stabilize training without sacrificing convergence speed. Zhu et al. [50] analyze the heavy-tailed nature of stochastic gradients in deep learning and derive generalization bounds that explicitly depend on the tail index $\alpha$, offering an explanation for the empirical robustness of SGD under non-Gaussian noise. Finally, Liu and He [29] extend these ideas to online convex optimization with heavy-tailed stochastic gradients, providing convergence guarantees and adaptive algorithms that remain stable even when gradient moments of order two do not exist. In our work, we adopt the empirical reality of heavy-tailed gradient noise as a basic modeling premise.
Finally, structured state-space models achieve long-range dependencies via implicit recurrence or convolutional kernels [14, 15]. Our focus remains on classical recurrent systems trained by BPTT: we deliver measurable per-lag detection thresholds and learnability windows backed by heavy-tailed gradient statistics.
3 Backpropagation through time
Training RNNs follows the same fundamental principle as feedforward models: parameter updates are obtained through SGD or variations of thereof [40]. Given trainable parameters $\theta$ and learning rate $\mu$, the update at iteration $r$ is
$\theta_{r+1}\;=\;\theta_{r}-\mu\,\nabla_{\theta}\mathcal{L}(\theta_{r}),\qquad\mathcal{L}\;=\;\sum_{t=1}^{T}\mathcal{E}_{t},$ (1)
where $\mathcal{E}_{t}$ denotes the instantaneous loss at time $t$ within a sequence of length $T$. In recurrent architectures, however, computing the gradient $\nabla_{\theta}\mathcal{L}$ requires unrolling the network dynamics through time and accounting for how earlier states influence later losses–a process known as Backpropagation Through Time.
Let $s_{t}$ denote the recurrent state. The total gradient of the loss with respect to the parameters can be written as
$\nabla_{\theta}\mathcal{L}\;=\;\sum_{t=1}^{T}\frac{\partial\mathcal{E}_{t}}{\partial\theta}\;=\;\sum_{t=1}^{T}\frac{\partial\mathcal{E}_{t}}{\partial s_{t}}\sum_{\ell=1}^{t}\frac{\partial s_{t}}{\partial s_{\ell}}\frac{\partial s_{\ell}}{\partial\theta}.$ (2)
The outer sum aggregates contributions from all time steps, while the inner chain rule expresses how parameter perturbations at earlier times influence the current loss through the recurrent dynamics. To make this dependence explicit, define $J_{j}\;=\;\frac{\partial s_{j}}{\partial s_{j-1}}$ and $B_{\ell}(\theta)\;=\;\frac{\partial s_{\ell}}{\partial\theta}$. Here $J_{j}$ is the state Jacobian, which quantifies how the state evolves in response to infinitesimal perturbations of the previous state, and $B_{\ell}(\theta)$ is the parameter-state Jacobian, measuring the instantaneous sensitivity of the state to the parameters at step $\ell$. Substituting these definitions into Eq. (2), the gradient contribution of a specific loss term $\mathcal{E}_{t}$ becomes
$\frac{\partial\mathcal{E}_{t}}{\partial\theta}\;=\;\delta_{t}^{\top}\sum_{\ell=1}^{t}\mathcal{M}_{t,\ell}\,B_{\ell}(\theta),\qquad\mathcal{M}_{t,\ell}\;:=\;\prod_{j=\ell+1}^{t}J_{j},$ (3)
where $\delta_{t}=\partial\mathcal{E}_{t}/\partial s_{t}$ denotes the local loss gradient at time $t$. The matrix product $\mathcal{M}_{t,\ell}$, often referred to as the state-transition Jacobian product, transports this signal backward through the sequence, modulating both its magnitude and direction as it interacts with the intermediate state Jacobians $J_{j}$.
4 Effective learning rates for LSTM and GRU
This section derives, for LSTM and GRU, the per-neuron, per-lag effective learning rates obtained via a first-order expansion of Jacobian products first introduced in [30] and briefly described in Appendix A. Together with the BPTT decomposition in Sec. 3, these rates connect state-space multiplicative geometry (gates) with parameter-space learning dynamics, and form the basis of the learnability analysis in Sec. 5.
4.1 General notation and conventions
Let $x_{t}\in\mathbb{R}^{D}$ be the input at step $t$, $h_{t}\in\mathbb{R}^{H}$ the hidden state, and (for LSTM) $c_{t}\in\mathbb{R}^{H}$ the cell state. For any $v\in\mathbb{R}^{H}$, let $D(v):=\mathrm{diag}(v)\in\mathbb{R}^{H\times H}$. Hadamard (elementwise) product is $\odot$. The logistic and hyperbolic tangent are elementwise: $\sigma(\cdot)$ and $\tanh(\cdot)$. We use the diagonal “slope” matrices
$S^{\sigma}(u):=D(\sigma^{\prime}(u)),\qquad S^{\tanh}(u):=D\!\big(1-\tanh^{2}(u)\big).$
When convenient, $D(h_{t-1})$ is abbreviated by $D_{h,t-1}$. Weight matrices have shapes $W_{\bullet}\in\mathbb{R}^{H\times D}$, $U_{\bullet}\in\mathbb{R}^{H\times H}$, and $b_{\bullet}\in\mathbb{R}^{H}$.
Throughout, one-step Jacobians are derivatives with respect to the previous state: for LSTM we stack $s_{t}=[h_{t};c_{t}]\in\mathbb{R}^{2H}$ and write $J_{t}=\partial s_{t}/\partial s_{t-1}\in\mathbb{R}^{2H\times 2H}$; for GRU, $J_{t}=\partial h_{t}/\partial h_{t-1}\in\mathbb{R}^{H\times H}$.
For a square $A\in\mathbb{R}^{n\times n}$, $\mathrm{diagvec}(A)\in\mathbb{R}^{n}$ is the vector formed by the diagonal of $A$. In LSTM, when $\mathcal{M}_{t,\ell}=\prod_{j=\ell+1}^{t}J_{j}$ is written in $2H\times 2H$ block form for $s=[h;c]$, the notation $\cdot\,]_{h,c}\in\mathbb{R}^{H\times H}$ refers to the top-right block mapping $c_{\ell}\mapsto h_{t}$.
4.2 LSTM: one-step Jacobian and effective learning rates
Dynamics.
An LSTM maintains $(h_{t},c_{t})$ and computes
$a^{i}_{t}$ $=W_{i}x_{t}+U_{i}h_{t-1}+b_{i},$ $i_{t}$ $=\sigma(a^{i}_{t}),$ (input gate) (4)
$a^{f}_{t}$ $=W_{f}x_{t}+U_{f}h_{t-1}+b_{f},$ $f_{t}$ $=\sigma(a^{f}_{t}),$ (forget/retention) (5)
$a^{o}_{t}$ $=W_{o}x_{t}+U_{o}h_{t-1}+b_{o},$ $o_{t}$ $=\sigma(a^{o}_{t}),$ (output/expression) (6)
$a^{g}_{t}$ $=W_{g}x_{t}+U_{g}h_{t-1}+b_{g},$ $g_{t}$ $=\tanh(a^{g}_{t}),$ (cell candidate) (7)
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot g_{t},\qquad h_{t}\;=\;o_{t}\odot\tanh(c_{t}).$ (8)
Define the diagonal gate matrices and slopes
$F_{t}=D(f_{t}),\ I_{t}=D(i_{t}),\ O_{t}=D(o_{t}),\ G_{t}=D(g_{t}),\quad S^{i}_{t}=S^{\sigma}(a^{i}_{t}),\ S^{f}_{t}=S^{\sigma}(a^{f}_{t}),\ S^{o}_{t}=S^{\sigma}(a^{o}_{t}),\ S^{g}_{t}=S^{\tanh}(a^{g}_{t}),$
and cell-expression factors
$S_{t}:=S^{\tanh}(c_{t})=D\!\big(1-\tanh^{2}(c_{t})\big),\quad H_{t}:=D\!\big(\tanh(c_{t})\big),\quad E_{t}:=D\!\big(o_{t}\odot(1-\tanh^{2}(c_{t}))\big)=O_{t}S_{t}.$
One-step Jacobian.
With $s_{t}=[h_{t};c_{t}]$, the Jacobian blocks are
$\frac{\partial c_{t}}{\partial c_{t-1}}$ $=F_{t},$ $\frac{\partial c_{t}}{\partial h_{t-1}}$ $=\underbrace{D(c_{t-1})\,S^{f}_{t}U_{f}}_{\text{via forget}}+\underbrace{I_{t}S^{g}_{t}U_{g}}_{\text{via candidate}}+\underbrace{G_{t}S^{i}_{t}U_{i}}_{\text{via input}}=:C_{t}^{(h)},$ (9)
$\frac{\partial h_{t}}{\partial c_{t-1}}$ $=E_{t}F_{t},$ $\frac{\partial h_{t}}{\partial h_{t-1}}$ $=H_{t}S^{o}_{t}U_{o}+E_{t}C_{t}^{(h)}.$ (10)
Collecting terms,
$J_{t}=\begin{bmatrix}H_{t}S^{o}_{t}U_{o}+E_{t}C_{t}^{(h)}&E_{t}F_{t}\\ C_{t}^{(h)}&F_{t}\end{bmatrix}\in\mathbb{R}^{2H\times 2H}.$ (11)
First-order expansion of the transport.
Let $\mathcal{M}_{t,\ell}=\prod_{j=\ell+1}^{t}J_{j}$. Decompose $J_{t}=T_{t}+R_{t}$ with
$T_{t}=\begin{bmatrix}0&E_{t}F_{t}\\ 0&F_{t}\end{bmatrix},\qquad R_{t}:=J_{t}-T_{t},$
so $T_{t}$ contains retention ($F_{t}$) and expression ($E_{t}$) along the cell path, while $R_{t}$ collects recurrently mixed corrections. By the first-order product rule,
$\mathcal{M}_{t,\ell}\ \approx\ T_{t,\ell}\ +\ \sum_{p=\ell+1}^{t}T_{t,p+1}\,R_{p}\,T_{p-1,\ell},\qquad T_{a:b}:=\prod_{j=b+1}^{a}T_{j},\;\;T_{b:b}:=I.$ (12)
Since the loss depends on $h_{t}$ (Sec. 3), only the top-right block of $\mathcal{M}_{t,\ell}$ (mapping $c_{\ell}\mapsto h_{t}$) contributes directly to the gradient transport.
Zeroth-order contributions.
Because $T_{t}$ is block upper-triangular with a zero $h\!\to\!h$ block, the $(h,c)$ block of the transport $T_{t,\ell}:=\prod_{j=\ell+1}^{t}T_{j}$ reduces to a single surviving path in which the signal remains in the cell state until time $t$:
$T_{t,\ell}\,]_{h,c}=E_{t}\,\Phi_{t:\ell},\qquad\Phi_{a:b}:=\prod_{j=b+1}^{a}F_{j},\;\;\Phi_{b:b}=I.$
In particular, the $c\!\to\!h$ transition can occur only at the final time $t$. Extracting the diagonal gives the neuron-wise rates
$\gamma^{(0)}_{t,\ell}:=\mathrm{diagvec}\!\left(E_{t}\,\Phi_{t:\ell}\right)\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=e_{t,q}\!\prod_{j=\ell+1}^{t}f_{j,q},\ \ e_{t,q}=o_{t,q}\big(1-\tanh^{2}(c_{t,q})\big).$ (13)
First-order diagonal correction.
From the first-order product expansion (12), the contribution of a single recurrent mixing insertion to the top-right block of the transport is
$\gamma^{(1)}_{t,\ell}\ :=\ \sum_{p=\ell+1}^{t}\mathrm{diagvec}\!\Big([\,T_{t,p}R_{p}T_{p-1,\ell}\,]_{h,c}\Big)\in\mathbb{R}^{H},$
where $\cdot\,]_{h,c}$ denotes the $H\times H$ block mapping $c_{\ell}\mapsto h_{t}$. Unlike the zeroth-order case, the presence of the recurrent correction $R_{p}$ reintroduces $h\!\to\!h$ transport, allowing a single mixing event to occur at any intermediate time $p\in\{\ell{+}1,\dots,t\}$. This term therefore collects the diagonal components of the first-order Fréchet expansion in Appendix A, corresponding to neuron-wise self-couplings induced by a single insertion of the recurrent mixing blocks. Off-diagonal entries encode cross-neuron interactions and are intentionally discarded, as they do not admit a neuron-wise time-scale interpretation.
LSTM effective learning rates.
With global learning rate $\mu>0$, we define the per-lag, per-neuron effective learning rates
$\mu_{t,\ell}\ :=\ \mu\big(\gamma^{(0)}_{t,\ell}+\gamma^{(1)}_{t,\ell}\big)\in\mathbb{R}^{H},\qquad\mu^{(q)}_{t,\ell}=\mu\big(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\big)$. (14)
These quantities act as scalar multipliers that modulate the contribution of gradients originating at lag $(t{-}\ell)$ to parameter updates, providing a neuron-wise characterization of the effective learning dynamics induced by the LSTM gates.
4.3 GRU: one-step Jacobian and effective learning rates
Dynamics.
A GRU updates
$a^{z}_{t}$ $=W_{z}x_{t}+U_{z}h_{t-1}+b_{z},$ $z_{t}$ $=\sigma(a^{z}_{t})$ (update/leak) (15)
$a^{r}_{t}$ $=W_{r}x_{t}+U_{r}h_{t-1}+b_{r},$ $r_{t}$ $=\sigma(a^{r}_{t})$ (reset/filtering) (16)
$a^{g}_{t}$ $=W_{h}x_{t}+U_{h}(r_{t}\odot h_{t-1})+b_{h},$ $g_{t}$ $=\tanh(a^{g}_{t})$ (candidate) (17)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot g_{t}.$ (18)
Let $Z_{t}=D(z_{t})$, $R_{t}=D(r_{t})$, $G_{t}=D(g_{t})$, $D_{h,t-1}=D(h_{t-1})$ and $S^{z}_{t}=S^{\sigma}(a^{z}_{t})$, $S^{r}_{t}=S^{\sigma}(a^{r}_{t})$, $S^{g}_{t}=S^{\tanh}(a^{g}_{t})$.
One-step Jacobian.
Differentiating the leak and update paths gives
$\frac{\partial}{\partial h_{t-1}}\!\big[(1-z_{t})\odot h_{t-1}\big]$ $=(I-Z_{t})\;-\;D_{h,t-1}\,S^{z}_{t}U_{z},$ (19)
$\frac{\partial g_{t}}{\partial h_{t-1}}$ $=S^{g}_{t}U_{h}\left(R_{t}+D_{h,t-1}S^{r}_{t}U_{r}\right),$ (20)
$\frac{\partial}{\partial h_{t-1}}\!\big[z_{t}\odot g_{t}\big]$ $=G_{t}S^{z}_{t}U_{z}\;+\;Z_{t}\,\frac{\partial g_{t}}{\partial h_{t-1}}.$ (21)
Summing contributions,
$J_{t}=(I-Z_{t})+(G_{t}-D_{h,t-1})\,S^{z}_{t}U_{z}+Z_{t}S^{g}_{t}U_{h}R_{t}+Z_{t}S^{g}_{t}U_{h}D_{h,t-1}S^{r}_{t}U_{r}.$ (22)
First-order expansion of the transport.
Write the Jacobian as $J_{t}=T_{t}+\mathcal{R}_{t}$, where
$T_{t}:=I-Z_{t}$
is the diagonal retention operator induced by the update gate, and $\mathcal{R}_{t}$ collects all remaining recurrent mixing terms. For the transport $\mathcal{M}_{t,\ell}=\prod_{j=\ell+1}^{t}J_{j}$, a first-order expansion yields
$\mathcal{M}_{t,\ell}\ \approx\ \Phi_{t:\ell}\ +\ \sum_{p=\ell+1}^{t}\Phi_{t:p}\,\mathcal{R}_{p}\,\Phi_{p-1:\ell},\qquad\Phi_{t:\ell}:=\prod_{j=\ell+1}^{t}(I-Z_{j}),\ \ \Phi_{\ell:\ell}=I.$ (23)
Zeroth-order contributions.
The diagonal product $\Phi_{t:\ell}$ yields the primary leak/retention envelope
$\gamma^{(0)}_{t,\ell}=\mathrm{diagvec}(\Phi_{t:\ell})\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-z_{j,q}),$ (24)
which governs the zeroth-order neuron-wise time scales associated with the update gate.
In addition, the multiplicative action of the reset gate along the candidate pathway induces further diagonal attenuations that are useful to track explicitly. We therefore define the diagonal pathwise envelopes
$\rho^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}r_{j,q},\qquad\eta^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-z_{j,q})\,r_{j,q},$ (25)
which capture pure reset attenuation and mixed update-reset attenuation, respectively. These quantities isolate the shrinkage associated with reset-controlled chains and are reported as complementary diagonal envelopes, rather than as additional terms in the strict $T_{t}+\mathcal{R}_{t}$ product expansion.
First-order diagonal correction.
The first-order diagonal correction associated with recurrent mixing is obtained by extracting the diagonal of the first–order term in (23):
$\gamma^{(1)}_{t,\ell}\ :=\ \sum_{p=\ell+1}^{t}\mathrm{diagvec}\!\big(\Phi_{t:p}\,\mathcal{R}_{p}\,\Phi_{p-1:\ell}\big)\in\mathbb{R}^{H}.$
This term collects the diagonal components arising from a single insertion of the recurrent mixing operator into an otherwise gate-diagonal transport. As in the LSTM case, off-diagonal components encode cross-neuron interactions and are intentionally discarded in order to preserve a neuron-wise time-scale interpretation.
GRU effective learning rates.
With global learning rate $\mu>0$, we define the per-lag, per-neuron effective learning rates as
$\mu^{(q)}_{t,\ell}\ :=\ \mu\Big(\gamma^{(0,q)}_{t,\ell}+\rho^{(0,q)}_{t,\ell}+\eta^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\Big),\qquad q=1,\dots,H.$ (26)
These quantities provide a diagonal, neuron-wise summary of how update and reset gates jointly shape the effective learning dynamics across temporal lags.
4.4 Considerations in terms of optimization dynamics
The effective learning rates $\mu_{t,\ell}=(\mu^{(1)}_{t,\ell},\ldots,\mu^{(H)}_{t,\ell})$ in (14) and (26) quantify how gates re-weight BPTT contributions at lag $(t{-}\ell)$ on a per-neuron basis, even under a fixed global learning rate $\mu$ in SGD (1). Zeroth-order terms produce diagonal contributions—retention/expression in LSTM, leak/reset/mixed in GRU—while first-order terms introduce anisotropy: the recurrent matrices $U_{\bullet}$ modulated by gate slopes mix coordinates and steer updates into privileged subspaces. As a result, gated RNNs act as implicit multi-rate optimizers: they set heterogeneous, lag-dependent learning rates across neurons and selectively bias update directions. This interpretation of gates as multi-rate optimizers was initially discovered in [30] for the gated RNN models in Appendix B.
5 Learnability from effective learning rates under heavy-tailed gradient noise
[…]
5.1 Setup and link to BPTT
We begin by isolating the portion of the BPTT gradient that reflects the influence of past states. Fix a lag $\ell\!\geq\!1$ and a time index $t$ with $t-\ell\!\geq\!1$. Let $s_{t}$ denote the recurrent state used by BPTT (Sec. 3); for LSTM, $s_{t}=[h_{t};c_{t}]$, whereas for GRU and gated RNNs in Appendix B, $s_{t}=h_{t}$.
Recalling the one-step decomposition in Eq. (3), we isolate the contribution of a past state $s_{t-\ell}$ to the current gradient. Define
$g_{t,\ell}(\theta)\;:=\;\delta_{t}^{\top}\mathcal{M}_{t,t-\ell}\,B_{t-\ell}(\theta)\;\in\;\mathbb{R}^{1\times P},$
where $\delta_{t}=\partial\mathcal{L}_{t}/\partial s_{t}$ is the local loss gradient, $\mathcal{M}_{t,t-\ell}$ is the $\ell$-step Jacobian product
$\mathcal{M}_{t,t-\ell}:=\prod_{k=t-\ell+1}^{t}\frac{\partial s_{k}}{\partial s_{k-1}},$
and $B_{t-\ell}(\theta)=\partial s_{t-\ell}/\partial\theta$ is the parameter Jacobian. The vector $g_{t,\ell}(\theta)$ represents the lag-$\ell$ contribution of the state $s_{t-\ell}$ to the BPTT gradient in parameter space.
Following the first-order approximation introduced in Appendix A, we approximate the Jacobian product by a diagonal operator collecting neuronwise effective learning rates:
$\mathcal{M}_{t,t-\ell}\;\approx\;D_{t,\ell}\;:=\;\mathrm{diag}\!\big(\mu^{(1)}_{t,\ell},\dots,\mu^{(H)}_{t,\ell}\big)$,
where $\mu^{(q)}_{t,\ell}\geq 0$ quantifies the cumulative modulation induced by gating along the $\ell$-step path associated with neuron $q$.
Decomposing the parameter Jacobian into neuronwise blocks
$B_{t-\ell}(\theta)=[B^{(1)}_{t-\ell}\ \cdots\ B^{(H)}_{t-\ell}],\qquad B^{(q)}_{t-\ell}\in\mathbb{R}^{1\times P},$
the lag-$\ell$ gradient contribution admits the following vector-valued first-order approximation:
$g_{t,\ell}(\theta)\;\approx\;\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\delta_{t}^{(q)}\,B^{(q)}_{t-\ell}(\theta),$ (27)
where $\delta_{t}^{(q)}$ denotes the $q$-th component of the local loss gradient at time $t$.
Equation (27) makes explicit that the lag-$\ell$ contribution to the parameter gradient is a vector obtained by summing neuronwise parameter sensitivities, each modulated by a corresponding effective learning rate. Thus, BPTT transports gradients along multiplicative chains of effective learning rate values, while the geometric structure of the parameter Jacobian determines how this transported signal is distributed across parameters.
5.2 Binary detection and matched statistic
We formalize finite-horizon learnability by asking when information about a past state $s_{t-\ell}$ remains statistically present in the stochastic gradient at time $t$. As shown in Sec. 5.1, the lag-$\ell$ contribution of $s_{t-\ell}$ to the BPTT gradient is the vector $g_{t,\ell}(\theta)\in\mathbb{R}^{1\times P}$ defined in Eq. (27).
[…]
We therefore compress the vector-valued quantity $g_{t,\ell}(\theta)$ into a scalar by applying a fixed linear readout. Let $w\in\mathbb{R}^{P}$ be a random unit-norm vector, drawn independently of the data and fixed throughout training and across all model instances. This projection defines a one-dimensional observable that can be interpreted as a random but unbiased probe of the gradient geometry.
Applying this readout to $g_{t,\ell}(\theta)$ and using the first-order approximation Eq. (27), we obtain the scalar neuronwise alignment variables
$\zeta^{(q)}_{t,\ell}:=\delta_{t}^{(q)}\,\langle B^{(q)}_{t-\ell}(\theta),\,w\rangle,$ (28)
so that the lag-$\ell$ contribution reduces to the weighted sum $\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\zeta^{(q)}_{t,\ell}$.
Because the same realization of $w$ is used for all models and training runs, differences in detectability across architectures reflect differences in temporal credit assignment and gradient transport, rather than artifacts of the projection itself. Moreover, since $w$ is drawn uniformly from the unit sphere, its distribution is rotationally invariant. As a result, the projection $\langle B^{(q)}_{t-\ell}(\theta),\,w\rangle$ provides an unbiased random probe of the typical magnitude and relative alignment of the neuronwise parameter-sensitivity directions, without privileging any fixed parameter direction.
The scalar variables $\zeta^{(q)}_{t,\ell}$ quantify how the transported gradient components associated with neuron $q$ align, through the parameter Jacobian, with the fixed readout direction $w$. The effective learning rates $\mu^{(q)}_{t,\ell}$ weight these alignment terms and thus govern how strongly each lag influences parameter updates.
Framed this way, the question of whether RNNs can use BPTT to exploit a dependency at lag $\ell$ becomes a question of detectability: is the signal induced by $s_{t-\ell}$ large enough, relative to gradient noise, to remain statistically distinguishable? This viewpoint allows us to apply information-theoretic tools to characterize when past states remain recoverable during training.
Define the expected alignment $m_{q}(\ell)=\mathbb{E}[\zeta^{(q)}_{t,\ell}$] and construct a matched statistic that aggregates lag-$\ell$ evidence across neurons:
$T_{t,\ell}=\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\mathrm{sgn}\!\big(m_{q}(\ell)\big)\,\zeta^{(q)}_{t,\ell}$. (29)
The factor $\mathrm{sgn}(m_{q}(\ell))$ flips each coordinate so that its contribution is oriented along the direction of its expected alignment. Consequently, under the data distribution, $\mathrm{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}$ has nonnegative mean, and every term in the sum contributes constructively. [^2] Formally, $\mathbb{E}\big[\mathrm{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}\big]=|m_{q}(\ell)|$ by definition of $m_{q}(\ell)$. Thus the expected signal contained in $T_{t,\ell}$ is always nonnegative, irrespective of the sign pattern of the raw alignments $m_{q}(\ell)$. The sign $\mathrm{sgn}(m_{q}(\ell))$ should be understood as a population-level orientation defining an ideal matched statistic; in practice, detectability depends only on the magnitude of the mean shift and is assessed from the empirical average of $T_{t,\ell}$, without requiring prior knowledge of this sign.
Taking expectations over the randomness of the data yields
$\mathbb{E}[T_{t,\ell}]=\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,|m_{q}(\ell)|=\overline{m}_{\mu}(\ell)\,f(\ell),$ (30)
where
$\overline{m}_{\mu}(\ell):=\frac{\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,|m_{q}(\ell)|}{\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}}$ (31)
is a weighted average of the neuronwise alignment magnitudes $|m_{q}(\ell)|$, with weights given by the corresponding effective learning rates.
This factorization separates two complementary aspects of temporal credit assignment: (i) $\overline{m}_{\mu}(\ell)$ captures the average informational alignment between the gradient signal and the parameter-sensitivity directions at lag $\ell$, and (ii) $f(\ell)$ quantifies the aggregate gain describing how much of that information survives recurrent transport through gates and state dynamics. Together, they determine the expected strength of the lagged contribution to the overall gradient, that is, how detectable past dependencies remain after being filtered by the network’s temporal mechanisms.
Averaging $T_{t,\ell}$ over $N$ independent training sequences yields $\widehat{T}_{N}(\ell)=\tfrac{1}{N}\sum_{n=1}^{N}T^{(n)}_{t,\ell}$, the empirical matched statistic that forms the basis for the finite-sample analysis in the next section.
5.3 Finite-sample analysis
We derive the finite-sample requirements for detecting a lag-$\ell$ dependency from noisy gradient information.
5.3.1 Statistical model
[…]
We extend the zero-mean formulation of [43] to a two-parameter location family, allowing small shifts in the mean corresponding to the presence or absence of a detectable signal. At lag $\ell$, the averaged matched statistic $\widehat{T}_{N}(\ell)$ over $N$ independent training sequences is modeled as
$\widehat{T}_{N}(\ell)\;\big|\;B\;\sim\;\mathcal{S}\alpha\mathcal{S}\!\Big(\mu_{\text{out}},\;\sigma_{\alpha}(\ell)/N^{1-1/\alpha}\Big),\qquad\mu_{\text{out}}\!\in\!\left\{+\tfrac{1}{2}\Delta(\ell),\;-\tfrac{1}{2}\Delta(\ell)\right\},$ (32)
where $\mu_{\text{out}}$ is the location parameter, the mean separation $\Delta(\ell)=\overline{m}_{\mu}(\ell)f(\ell)$ quantifies the strength of the lag-$\ell$ signal, $\sigma_{\alpha}(\ell)$ acts as a noise-scale proxy, and $B\in\{0,1\}$ describes the non-detection/detection outcome.
For analytical tractability, we assume bounded alignment and noise scales, $c_{m}\leq m_{\mu}(\ell)\leq C_{m}$ and $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$, which guarantee well-defined sample-complexity constants without affecting the asymptotic exponents. The factor $N^{1-1/\alpha}$ encodes the slow concentration typical of $\alpha$-stable averages, and a justification for this choice is discussed in Appendix C.
5.3.2 Information-theoretic bounds
[…]
$P_{\text{det}}=\mathcal{S}\alpha\mathcal{S}\!\Big({+}\tfrac{1}{2}\Delta(\ell),\,\sigma_{\alpha}(\ell)/N^{\,1-1/\alpha}\Big),\qquad P_{\text{non}}=\mathcal{S}\alpha\mathcal{S}\!\Big({-}\tfrac{1}{2}\Delta(\ell),\,\sigma_{\alpha}(\ell)/N^{\,1-1/\alpha}\Big),$
where $\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell)$ denotes the mean separation.
For $\alpha>1$, symmetric $\alpha$–stable location families admit local asymptotic normality (LAN) with finite Fisher information for the location parameter [20, 26, 35], sufficient for information-theoretic bounds. In the present setting, the observation $\widehat{T}_{N}(\ell)$ follows a triangular array of location models [45] whose scale shrinks as $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$. For such families, LAN applies with normalization rate $r_{N}=1/s_{N}$. As a consequence, the log-likelihood ratio between $P_{\text{det}}$ and $P_{\text{non}}$ admits a quadratic expansion for small location shifts, yielding the lower bound for the Kullback-Leibler (KL) divergence
$D_{\mathrm{KL}}\!\left(P_{\text{det}}\|P_{\text{non}}\right)\;\geq\;c_{\alpha}\,\frac{N^{\,2(1-1/\alpha)}\,\Delta(\ell)^{2}}{\sigma_{\alpha}(\ell)^{2}},$ (33)
where $c_{\alpha}>0$ depends only on $\alpha$. A derivation of this bound for shrinking-scale $\alpha$-stable location models is provided in Appendix D.
Combining (33) with standard relations between KL divergence and mutual information [11] yields
$I\!\left(B;\widehat{T}_{N}(\ell)\right)\;\geq\;\min\!\left\{\log 2,\;c_{\alpha}\,\frac{N^{\,2(1-1/\alpha)}}{\sigma_{\alpha}(\ell)^{2}}\,\overline{m}_{\mu}(\ell)^{2}\,f(\ell)^{2}\right\}.$ (34)
Applying Fano’s inequality [11] to the binary detection problem yields the minimal number of independent sequences $N$ required to achieve a target error probability $P_{e}\leq\epsilon$:
$N\;\geq\;\left(\frac{\sigma_{\alpha}(\ell)}{\sqrt{c_{\alpha}}\,\overline{m}_{\mu}(\ell)f(\ell)}\right)^{\!\frac{\alpha}{\alpha-1}}\!\!\Big(\log\tfrac{1}{2\epsilon}\Big)^{\frac{\alpha}{2(\alpha-1)}}.$ (35)
This expression links sample complexity to the noise scale $\sigma_{\alpha}(\ell)$, the average alignment $\overline{m}_{\mu}(\ell)$, and the envelope $f(\ell)$. In contrast to the Gaussian case, the exponent $\alpha/(\alpha-1)>2$ for $\alpha<2$, reflecting the slower statistical concentration induced by heavy-tailed gradient noise. Full derivations of Eqs. (34) and (35) are given in Appendix E.
5.4 Learnability window and scaling laws
Throughout this section we adopt the corrected concentration exponent
$\kappa_{\alpha}\;:=\;\frac{\alpha}{\alpha-1},\qquad\alpha>1,$ (36)
which governs the finite-sample detectability of lagged dependencies under $\alpha$-stable gradient noise when statistics are averaged over independent sequences. Equivalently, the noise floor of the empirical matched statistic decays as $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$.
Starting from the finite-sample condition in Eq. (35), we can invert the relation to express the minimal effective learning rate mass required for reliable detection at a given sample size $N$. Rearranging terms [^3] To obtain Eq. (37) from Eq. (35), one takes the $\kappa_{\alpha}$th root of the bound, isolates the factor $1/f(\ell)$, and inverts the (positive) inequality. This yields the minimal effective learning rate mass required for detectability at lag $\ell$. yields the per-lag detectability threshold
$\varepsilon_{\mathrm{th}}(\ell)=\frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)}\cdot\frac{1}{\sqrt{c_{\alpha}}}\,\sqrt{\log\!\Big(\tfrac{1}{2\epsilon}\Big)}.$ (37)
This threshold compactly summarizes the interplay of noise scale $\sigma_{\alpha}(\ell)$, data size $N$, and alignment strength $\overline{m}_{\mu}(\ell)$: for a dependency at lag $\ell$ to be statistically detectable under heavy-tailed $\alpha$-stable noise, the envelope $f(\ell)$ must exceed $\varepsilon_{\mathrm{th}}(\ell)$. In this formulation, smaller $\alpha$ values (heavier tails) increase $\kappa_{\alpha}$, thereby slowing statistical concentration to $N^{1/\kappa_{\alpha}}$ and reducing the effective temporal range over which informative gradients can be recovered.
We can now define the learnability window.
Definition 5.1 (Learnability window under $\alpha$-stable noise).
The learnability window is
$\mathcal{H}_{N}=\sup\big\{\ell\geq 1:\,f(\ell)\geq\varepsilon_{\mathrm{th}}(\ell)\big\}.$ (38)
Intuitively, $\mathcal{H}_{N}$ is the largest lag for which the transported gradient retains a recoverable signal. Even if Jacobians are numerically stable, once $f(\ell)$ falls below $\varepsilon_{\mathrm{th}}(\ell)$, heavy-tailed fluctuations dominate and credit assignment becomes statistically infeasible.
From the Fano bound in Eq. (35), the minimal number of independent sequences sufficient to detect a lag-$\ell$ dependency can be written as
$N(\ell)\;:=\;\kappa_{\alpha,\epsilon}\,\left(\frac{\sigma_{\alpha}(\ell)}{\overline{m}_{\mu}(\ell)f(\ell)}\right)^{\!\kappa_{\alpha}},\qquad\kappa_{\alpha,\epsilon}\;:=\;\frac{1}{c_{\alpha}^{\kappa_{\alpha}/2}}\,\Big(\log\tfrac{1}{2\epsilon}\Big)^{\kappa_{\alpha}/2}.$ (39)
The factor $\kappa_{\alpha,\epsilon}$ depends only on the tail index $\alpha$ and the target error level $\epsilon$ (fixed across lags) and will be absorbed into constant factors in our asymptotic statements.
Here, we formalize two regularities that drive the scaling of $\mathcal{H}_{N}$. Proofs appear in Appendix F.
Lemma 5.1 (Monotonicity in the lag).
Fix $t$ and a neuron $q$. Assume gate activations lie in $[0,1]$ and activation derivatives are bounded in $[0,1]$. Then $\ell\mapsto\mu^{(q)}_{t,\ell}$ is nonincreasing and, consequently, $\ell_{1}\leq\ell_{2}\Rightarrow f(\ell_{2})\leq f(\ell_{1}).$
Lemma 5.2 (Sample-complexity scaling and window bounds).
Assume there exist constants $0<c_{\sigma}\leq C_{\sigma}$ and $0<c_{m}\leq C_{m}$ such that, over the lags of interest, $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$ and $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$. Then, with $N(\ell)$ defined in Eq. (39), there are constants $0<c_{\star}\leq C_{\star}$ for which
$c_{\star}\,f(\ell)^{-\kappa_{\alpha}}\ \leq\ N(\ell)\ \leq\ C_{\star}\,f(\ell)^{-\kappa_{\alpha}},\qquad c_{\star}:=\kappa_{\alpha,\epsilon}\Big(\tfrac{c_{\sigma}}{C_{m}}\Big)^{\!\kappa_{\alpha}},\quad C_{\star}:=\kappa_{\alpha,\epsilon}\Big(\tfrac{C_{\sigma}}{c_{m}}\Big)^{\!\kappa_{\alpha}}.$ (40)
Consequently, with $f^{\leftarrow}(y)=\sup\{\ell\geq 1:\ f(\ell)\geq y\}$ being the generalized inverse of $f(\ell)$,
$f^{\leftarrow}\!\Big(\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}\Big)\ \leq\ \mathcal{H}_{N}\ \leq\ f^{\leftarrow}\!\Big(\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\Big).$ (41)
The first lemma captures the monotone, multiplicative attenuation induced by gates; the second makes explicit that the per-lag sample complexity grows as the inverse $\kappa_{\alpha}$-power of the envelope $f(\ell)$. Together, they reduce the problem of characterizing $\mathcal{H}_{N}$ to inverting the decay profile of $f(\ell)$ at the level $N^{-1/\kappa_{\alpha}}$. In other words, once $\sigma_{\alpha}(\ell)$ and $\overline{m}_{\mu}(\ell)$ are bounded, the temporal reach of learnability is controlled by the speed with which $f(\ell)$ decays with $\ell$ and by the heavy-tailed concentration exponent $\kappa_{\alpha}$.
Representative scaling regimes.
In general, gated RNNs may exhibit more intricate behaviors for $f(\ell)$, including mixtures of regimes or multi-phase decay. To gain analytic insight, we focus on three prototypical asymptotic profiles that arise naturally from gate-induced multiplicative dynamics and that are observed in our experiments: logarithmic, polynomial, and exponential decay. These should be viewed as canonical regimes rather than an exhaustive catalogue; more complex envelopes can be analyzed by the same mechanism via Eq. (41).
For these three representative behaviors of $f(\ell)$ we obtain the following asymptotics (up to constants):
  1. (i)
    Logarithmic decay: if $f(\ell)\!\asymp\!c/\log(1+\ell)$, then $N(\ell)\!\asymp\![\log(1+\ell)]^{\kappa_{\alpha}}$ and $\mathcal{H}_{N}\!\asymp\!\exp(\kappa N^{1/\kappa_{\alpha}})-1$.
  2. (ii)
    Polynomial decay: if $f(\ell)\!\asymp\!c\,\ell^{-\beta}$, then $N(\ell)\!\asymp\!\ell^{\kappa_{\alpha}\beta}$ and $\mathcal{H}_{N}\!\asymp\!N^{1/(\kappa_{\alpha}\beta)}$.
  3. (iii)
    Exponential decay: if $f(\ell)\!\asymp\!c\,\lambda^{\ell}$ with $\lambda\!\in(0,1)$, then $N(\ell)\!\asymp\!\lambda^{-\kappa_{\alpha}\ell}$ and $\mathcal{H}_{N}\!\asymp\!(\log N)/[\kappa_{\alpha}\log(1/\lambda)]$.
In these expressions, the symbol $\asymp$ indicates “of the same asymptotic order”. These laws quantify how gating geometry and heavy-tailed noise jointly shape the temporal extent of learnability: exponential decay implies sharp forgetting and only logarithmic horizon growth; polynomial decay yields algebraic horizons; and logarithmic decay approaches a near-critical regime with rapidly expanding $\mathcal{H}_{N}$. In all cases, smaller $\alpha$ compresses $\mathcal{H}_{N}$ by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Full derivations and constants appear in Appendix G.
5.5 Theoretical implications
Combining the Fano-derived sample-complexity bound in Eq. (35) with the learnability window analysis of Sec. 5.4 yields several structural insights into how gating structure, heavy-tailed noise, and finite data interact.
(i) Gate geometry dominates data under heavy tails.
The detectability of a lag-$\ell$ dependency is determined by the signal strength $\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell)$, which (see Eq. (30)) grows linearly with the envelope $f(\ell)$. In the mutual information bound of Eq. (34), this signal strength appears squared as $\Delta(\ell)^{2}$, while the contribution of data enters through the factor $N^{2(1-1/\alpha)}=N^{2/\kappa_{\alpha}}$.
When $\alpha<2$, this growth is sublinear in $N$, reflecting the slow statistical concentration of heavy-tailed averages. Consequently, improving the gating geometry so that $f(\ell)$ decays more slowly produces a multiplicative increase in $\Delta(\ell)$, and hence a squared increase in the information term $\Delta(\ell)^{2}$. By contrast, increasing $N$ yields only a sublinear improvement whenever $\alpha<2$.
This imbalance is made explicit in the Fano bound (35), where the minimal sample size satisfies $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$. Thus, architectures whose gate spectra induce broader or more heterogeneous time scales, thereby maintaining larger values of $f(\ell)$ across lags, reduce the required sample size much more effectively than simply collecting more data. In this sense, under heavy-tailed noise, gate geometry dominates dataset size in controlling the learnability window $\mathcal{H}_{N}$.
(ii) The tail index $\alpha$ governs statistical efficiency.
The corrected concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$ quantifies how quickly noise averages out when statistics are aggregated across independent sequences. For Gaussian fluctuations ($\alpha=2$), one recovers the familiar concentration rate $N^{-1/2}$; for $\alpha<2$, the slower rate $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$ substantially weakens the effective signal-to-noise ratio. Thus, even with identical gating (and therefore identical $f(\ell)$), smaller $\alpha$ increases $\kappa_{\alpha}$, compresses $\mathcal{H}_{N}$, and raises the number of samples required for reliable detection.
(iii) Decay regime of $f(\ell)$.
The learnability window is determined by inverting the decay profile of $f(\ell)$ at the threshold level set by $N^{-1/\kappa_{\alpha}}$ (Lemma 5.2). Consequently: exponential decay yields logarithmic horizon growth (rapid forgetting); polynomial decay produces algebraic horizons (long memory); and logarithmic decay approaches a near-critical regime with quasi-exponential growth of $\mathcal{H}_{N}$. Across all regimes, reducing $\alpha$ uniformly tightens the horizon by slowing statistical concentration and shifting the detectability to shorter lags.
(iv) Architectural and training implications.
Architectures that promote slowly decaying envelopes are intrinsically more capable of learning long-range dependencies under heavy-tailed noise. Equally, training methods that reduce the impact of heavy-tailed fluctuations (e.g., clipping, normalization, or stabilizing adaptive optimizers) effectively increase $\alpha$, thereby reducing $\kappa_{\alpha}$ and improving sample efficiency without altering model size.
(v) Vanishing learnability as an information-theoretic obstruction.
The learnability window $\mathcal{H}_{N}$ may vanish identically for certain architectures, meaning that $\mathcal{H}_{N}=0$ for all finite $N$. In this case, the envelope $f(\ell)$ falls below the statistical detectability threshold for every $\ell\geq 1$, and the corresponding sample complexity $N(\ell)$ diverges. This does not reflect an optimization failure or insufficient data, but an information-theoretic obstruction: the transported gradients carry no statistically usable signal at nonzero lags. Equivalently, the KL divergence between the detection and non-detection models vanishes, so no estimator can reliably distinguish the two hypotheses regardless of sample size.
6 Empirical validation
We now validate the learnability theory on gated RNNs of variable complexity. Given the currently available computational resources, we are able to simulate the models defined in Appendix B: ConstGate (fixed scalar gate), SharedGate (learned scalar gate shared across units), and DiagGate (per-neuron gate). These models already show the qualitative regimes appearing in the theoretical analysis. Our goals are: (i) to empirically verify the monotone decay and scaling properties of the effective learning rate mass predicted by the theory; (ii) to compare the resulting learnability windows $\mathcal{H}_{N}$ across architectures with different gating structure; and (iii) to illustrate how the distribution of neuronwise time scales $\{\tau_{q}\}_{q}$ correlates with the empirical learnability window.
6.1 Experimental setup
Task and data.
We consider a synthetic regression task defined on input-output pairs $\{(x_{t},y_{t})\}_{t=1}^{T}$ with $x_{t}\in\mathbb{R}^{16}$ and scalar target $y_{t}\in\mathbb{R}$. The target is generated as a weighted sum of five delayed inputs,
$y_{t}=\sum_{k=1}^{5}c_{k}\,u^{\top}x_{t-\ell_{k}}+\varepsilon_{t},$ (42)
where $u\in\mathbb{R}^{16}$ is a random unit vector, the task lags are $\{\ell_{k}\}=\{32,\,64,\,128,\,256,\,512\}$, the mixing coefficients are $\{c_{k}\}=\{0.6,\,0.45,\,0.35,\,0.28,\,0.22\}$, and $\varepsilon_{t}$ is i.i.d. Gaussian noise with standard deviation $\sigma_{\text{noise}}=0.35$. These five lags induce informative dependencies across a broad temporal range.
To evaluate the model dynamics, all effective learning rate and sample-complexity quantities are computed over a separate diagnostic lag grid $\ell\in\{4,8,\ldots,128\}$ of $32$ evenly spaced values, chosen to probe short and intermediate temporal scales while avoiding numerical instabilities at very large lags. All training and diagnostic sequences are generated independently from the same process, with no reuse or overlap. This setup provides controlled long-range structure in the target while enabling a clean assessment of how each architecture’s envelope $f(\ell)$ shapes its empirical learnability window.
Architectures.
All recurrent models share the same backbone update
$h_{t}=(1-s_{t})\odot h_{t-1}+s_{t}\odot\tilde{h}_{t},\qquad\tilde{h}_{t}=\tanh(W_{h}x_{t}+U_{h}h_{t-1}+b_{h}),$
with hidden state $h_{t}\in\mathbb{R}^{H}$ and gate vector $s_{t}\in(0,1)^{H^{\prime}}$. The three RNN models differ only in the structure of the gate: (i) ConstGate uses a fixed scalar gate $s\in(0,1)$ shared across all units and all time steps (54); (ii) SharedGate uses a learned scalar gate $s_{t}\in(0,1)$ shared across units but varying with time (52); and (iii) DiagGate uses a learned per-neuron gate $s_{t}\in(0,1)^{H}$ updated coordinatewise (50).
All models use the same hidden dimension $H$, input dimension $D$, and a linear readout $y_{t}=w^{\top}h_{t}$ trained with mean-squared error loss. Recurrent weights $W_{h}$ and $U_{h}$ are initialized orthogonally and all biases are set to zero. For SharedGate and DiagGate, the gate parameters $(W_{s},U_{s},b_{s})$ are initialized such that the gate pre-activations are close to zero, producing initial gates $s_{t}\approx\tfrac{1}{2}$ and well-conditioned initial time scales. [^4] This follows the time-scale initialization heuristic proposed in [30].
Training protocol.
Each model is trained using mini-batch AdamW [31] with a constant base learning rate $\mu=10^{-3}$, weight decay $\lambda=10^{-4}$, and batch size $B=16$. Training uses $N_{\text{seq}}=8000$ independent sequences of length $T=1024$ (input dimension $D=16$, hidden size $H=64$) for $400$ epochs. Gradients are clipped to a maximum $\ell_{2}$ norm of $1.0$. No learning-rate schedule is applied. Unless otherwise noted, the same optimization hyperparameters are used for all architectures. After training, we freeze all network parameters and generate an independent diagnostic set of $N_{\text{diag}}=8000$ fresh sequences sampled from the same data-generating process, with no overlap in either time or random seed. All reported quantities are computed on this diagnostic set under the frozen post-training parameters.
Remark on the optimizer and gradient noise.
In our experiments we use AdamW rather than plain SGD, primarily for numerical stability and faster convergence. While the theoretical analysis in Sec. 5 is formulated in terms of empirical averages under $\alpha$-stable gradient noise, recent work has shown that heavy-tailed gradient fluctuations are not specific to vanilla SGD, but persist also for adaptive optimization methods such as Adam and AdaGrad. In particular, [7] provides both theoretical and empirical evidence that gradient noise remains heavy-tailed under adaptive optimizers, and that techniques such as gradient clipping improve their robustness without eliminating heavy-tail effects. These findings support the use of AdamW in our empirical study while remaining consistent with the heavy-tailed noise model underlying the learnability window analysis.
Effective learning rates.
[…]
Heavy-tailed gradient statistics and sample complexity.
To connect the empirical envelopes to the learnability theory, we estimate the heavy-tailed quantities appearing in the theoretical sample complexity expression. For each diagnostic lag $\ell$, we compute the empirical matched statistic $\widehat{T}_{t,\ell}$ defined in Sec. 5.2. From these samples, we estimate a global tail index $\hat{\alpha}$ and obtain a variance-free, robust estimate of the $\alpha$-stable scale parameter $\sigma_{\alpha}(\ell)$ via McCulloch’s quantile-based procedure [33]. We also estimate the alignment factor $\widehat{\overline{m}}_{\mu}(\ell)$ as the empirical mean of the matched statistic.
[…]
$\widehat{\mathcal{H}}_{N}=\max\{\ell:\widehat{N}(\ell)\leq N\},$ (43)
which serves as a data-driven surrogate for the theoretical horizon $\mathcal{H}_{N}$. Comparing $\widehat{\mathcal{H}}_{N}$ across architectures reveals how gate structure shapes temporal learnability in practice.
Time-scale extraction.
[…]
6.2 Empirical learnability windows and time-scale spectra
We now report results for the three gated RNNs in Appendix B. To keep figures readable, we group plots by quantity rather than by model: envelopes and scaling laws, learnability windows, and time-scale spectra.
Envelopes of the effective learning rates.
Figure 1 (left) shows the estimated envelopes $f(\ell)$ for ConstGate, SharedGate, and DiagGate. ConstGate and SharedGate exhibit rapid decay: their envelopes collapse to numerically negligible values by $\ell\approx 16$2020. In contrast, DiagGate retains substantial mass across the entire range, decaying much more slowly.
On a semi-logarithmic scale (Fig. 1, middle), the ConstGate and SharedGate curves become nearly straight lines with slopes corresponding to exponential decays $(1-s)^{\ell}$ with short time constants $\tau_{\text{env}}^{\text{const}}\approx 1.2$ and $\tau_{\text{env}}^{\text{shared}}\approx 1.8$. In agreement with the theory, an exponential fit $\hat{f}(\ell)\approx c\exp(-\ell/\tau_{\text{env}})$ achieves $r^{2}\approx 1$ for both models, while a power-law fit performs poorly.
For DiagGate, the semi-logarithmic plot reveals an almost flat line (very large $\tau_{\text{env}}\approx 46$), but the log-log plot (Fig. 1, right) is nearly linear with slope close to 1-1. A power-law fit $\hat{f}(\ell)\approx c\,\ell^{-\beta}$ yields an exponent $\beta\approx 1$ with $r^{2}\approx 0.98$, whereas the exponential fit is noticeably worse. Empirically, the aggregated effective learning rates of DiagGate behave like a polynomial function $\hat{f}(\ell)\propto\ell^{-1}$ over the probed range of lags, realizing the algebraic scaling regime.
(a) Linear scale
(b) Semi-log scale
(c) Log-log scale
Figure 1: Envelopes of the effective learning rates for ConstGate, SharedGate, and DiagGate models. (a) Linear view. (b) Semi-logarithmic view shows clean exponential decay for ConstGate and SharedGate. (c) Log-log view highlights the approximate polynomial behavior $\hat{f}(\ell)\propto\ell^{-1}$ for DiagGate.
Empirical learnability windows.
[…]
Time-scale spectra.
To connect the envelope geometry back to gating, we inspect the distribution of neuronwise time scales $\{\tau_{q}\}_{q}$ inferred from the effective learning rates. Figure 4 shows kernel density estimates of $\tau_{q}$ for each architecture. ConstGate and SharedGate both exhibit degenerate spectra: all units share the same time constant, yielding delta-like peaks at $\tau\approx 1.2$ and $\tau\approx 1.8$, respectively. DiagGate, in contrast, displays a broad, heavy-tailed distribution: most units have short time scales ($\tau\lesssim 3$), but a nontrivial fraction inhabit intermediate ranges ($\tau\approx 5$$10$), and at least one neuron develops a very long memory ($\tau\approx 60$).
This mixture of time scales explains the empirical power-law envelope of DiagGate: a superposition of exponentials with a broad spectrum of $\tau_{q}$ is well approximated, over a finite range of lags, by an algebraic kernel. Together, the results confirm the central message of Sec. 5: architectures with narrow gate spectra (ConstGate, SharedGate) lie firmly in the exponential regime and exhibit practically negligible learnability windows under heavy-tailed noise, whereas architectures that generate broad gate spectra (DiagGate) realize polynomial envelopes and substantially larger learnability windows at fixed sample budgets.
(a) ConstGate.
(b) SharedGate.
(c) DiagGate.
Figure 4: Distribution of neuronwise time scales $\{\tau_{q}\}$. (a) ConstGate exhibits a degenerate spectrum with a single characteristic time scale ($\tau\approx 1.2$), reflecting its fixed scalar gate. (b) SharedGate similarly collapses to one learned time constant ($\tau\approx 1.8$), showing no intra-network diversity. (c) DiagGate develops a broad and heavy-tailed spectrum: most units have short time scales ($\tau\lesssim 3$), but several occupy intermediate ranges and at least one neuron attains long memory ($\tau\approx 60$). This heterogeneous mixture of time scales explains the polynomial envelope $f(\ell)\propto\ell^{-1}$ observed for DiagGate and the substantially larger learnability window reported in Fig. 2.
7 Discussion and future directions
This work develops a quantitative framework linking gating geometry, heavy-tailed gradient noise, and the finite-horizon learnability of RNNs. A central message is that the effective learning rates $\mu_{t,\ell}$, rather than Jacobian stability in isolation, govern how much gradient signal survives across temporal lags. The decay profile of $f(\ell)$, together with the heavy-tailed index $\alpha$ through the concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$, determines the sample complexity curve $N(\ell)$ and thus the learnability window $\mathcal{H}_{N}$.
[…]
Several avenues for future work emerge naturally. First, extending the empirical evaluation to LSTM and GRU architectures, whose effective learning rates were derived in this paper, would test the generality of the framework. Although such simulations are computationally demanding and currently beyond our available hardware resources, they would provide valuable evidence on whether multi-gate architectures with heterogeneous gating mechanisms develop broad or even multimodal spectra of time scales, and whether their resulting learnability windows align with the theoretical predictions.
Second, our results point to a deeper open question: why do some architectures spontaneously develop broad time-scale spectra during training? The learnability framework introduced here characterizes the statistical consequences of these spectra, but not their origins. A complementary theory explaining the emergence (or collapse) of mixed time scales–potentially arising from the coupled dynamics of states, gates, and parameters–would provide a unified understanding of how long-range memory structures arise in practice.
Finally, an important direction concerns the interaction between heavy-tailed gradient statistics and modern optimization techniques. While the present theory is formulated in terms of empirical averages under $\alpha$-stable noise, understanding how adaptive optimizers, normalization, and gradient clipping modify (or preserve) heavy-tailed behavior remains an open problem. Clarifying when such mechanisms effectively increase the concentration rate, and when they merely truncate extremes without restoring Gaussian statistics, would further refine the practical scope of the learnability window framework.
Taken together, the learnability window offers a principled lens on the temporal limits of gated recurrent networks, clarifying when long-range dependencies are statistically detectable and how architectural design choices shape those limits. It provides a bridge between dynamical systems, heavy-tailed statistics, and sample complexity, revealing how gating structure–rather than data or optimization alone–fundamentally constrains the horizons over which RNNs can learn.
Appendix A Matrix product expansion via the Fréchet derivative formulation
This section summarizes the first-order expansion of a product of matrices with structured perturbations introduced in [30].
Definition A.1 (Fréchet differentiability [24, 17]).
Let $f:\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}$. We say that $f$ is Fréchet differentiable at $A\in\mathbb{C}^{n\times n}$ if there exists a bounded linear mapping $L_{f}(A,\cdot)$ such that
$\lim_{\|E\|\to 0}\frac{\|\,f(A+E)-f(A)-L_{f}(A,E)\,\|}{\|E\|}=0.$ (44)
If $f$ is Fréchet differentiable at $A$,
$f(A+E)\;=\;f(A)\;+\;L_{f}(A,E)\;+\;o(\|E\|).$ (45)
If $g$ and $h$ are Fréchet differentiable at $A$ and $f(X)=g(X)h(X)$, then
$L_{gh}(A,E)\;=\;L_{g}(A,E)\,h(A)\;+\;g(A)\,L_{h}(A,E).$ (46)
Consider
$F(\varepsilon)=\prod_{j=1}^{n}\big(A_{j}+\varepsilon B_{j}\big),$ (47)
with $A_{j},B_{j}\in\mathbb{C}^{d\times d}$. By recursive application of (46) one obtains
$L_{F_{n}}(0,E)\;=\;\sum_{i=1}^{n}\left(\prod_{j=1}^{i-1}A_{j}\right)\,B_{i}\,\left(\prod_{j=i+1}^{n}A_{j}\right),$ (48)
and the first-order expansion
$F(\varepsilon)\;=\;\left(\prod_{j=1}^{n}A_{j}\right)\;+\;\varepsilon\sum_{m=1}^{n}\left(\prod_{j=m+1}^{n}A_{j}\right)\,B_{m}\,\left(\prod_{j=1}^{m-1}A_{j}\right)\;+\;O(\varepsilon^{2}).$ (49)
Appendix B Diagonally gated RNNs: One-step Jacobians and effective learning rates
We introduce diagonally gated RNN models previously considered in [30] that we use throughout the paper: a per-neuron (diagonal) gate, a shared (global) gate, and a constant gate. All follow the same update template:
$h_{t}\;=\;(1-s_{t})\odot h_{t-1}\;+\;s_{t}\odot\tilde{h}_{t},\qquad\tilde{h}_{t}\;=\;\tanh(a^{h}_{t}),\;\;a^{h}_{t}=W_{h}x_{t}+U_{h}h_{t-1}+b_{h},$
but differ in how the gate $s_{t}$ is produced. We reuse the conventions of Sec. 4.
(a) Per-neuron (diagonal) gate $s_{t}\in(0,1)^{H}$ (“DiagGate”)
The gate is computed per coordinate:
$a^{s}_{t}\;=\;W_{s}x_{t}+U_{s}h_{t-1}+b_{s},\qquad s_{t}=\sigma(a^{s}_{t})\in(0,1)^{H}.$ (50)
Define $S_{t}:=D(s_{t})$ and $S^{s}_{t}:=S^{\sigma}(a^{s}_{t})$. The exact one-step Jacobian $J_{t}=\partial h_{t}/\partial h_{t-1}\in\mathbb{R}^{H\times H}$ is
$J_{t}\;=\;\underbrace{(I-S_{t})}_{\text{leak}}\;+\;\underbrace{\bigl(D(\tilde{h}_{t})-D(h_{t-1})\bigr)\,S^{s}_{t}U_{s}}_{\text{gate sensitivity}}\;+\;\underbrace{S_{t}\,S^{h}_{t}\,U_{h}}_{\text{candidate path}}.$ (51)
Here the leak term $(I-S_{t})$ is diagonal; the remaining two terms contain $U_{s}$ and $U_{h}$ and are generally full rank, mixing information across neurons.
Effective learning rates.
The zeroth-order envelope is given by
$\gamma^{(0,q)}_{t,l}=\prod_{j=l+1}^{t}(1-s_{j,q}),$
yielding per-neuron effective learning rates
$\mu^{(q)}_{t,l}=\mu\left(\gamma^{(0,q)}_{t,l}+\gamma^{(1,q)}_{t,l}\right)$,
where $\gamma^{(1,q)}_{t,l}$ arises from the diagonal of the gate-sensitivity and candidate-path corrections in Eq. (51).
(b) Shared (global) scalar gate $s_{t}\in(0,1)$ (“SharedGate”)
The gate is a single scalar per time step (and per sequence element in a batch):
$a^{s}_{t}\;=\;w_{s}^{\top}x_{t}+u_{s}^{\top}h_{t-1}+b_{s},\qquad s_{t}=\sigma(a^{s}_{t})\in(0,1).$ (52)
The update is $h_{t}=(1-s_{t})h_{t-1}+s_{t}\tilde{h}_{t}$ with $\tilde{h}_{t}=\tanh(a^{h}_{t})$. Using $\partial s_{t}/\partial h_{t-1}=s_{t}(1-s_{t})\,u_{s}$, the Jacobian is
$J_{t}\;=\;\underbrace{(1-s_{t})\,I}_{\text{leak}}\;+\;\underbrace{s_{t}\,S^{h}_{t}\,U_{h}}_{\text{candidate path}}\;+\;\underbrace{s_{t}(1-s_{t})\,(\tilde{h}_{t}-h_{t-1})\,u_{s}^{\top}}_{\text{rank-1 gate sensitivity}}.$ (53)
The last term is a rank-1 outer product and is the only source of cross-neuron coupling other than $U_{h}$.
Effective learning rates.
The zeroth-order envelope is
$\gamma^{(0)}_{t,l}=\prod_{j=l+1}^{t}(1-s_{j}),$
which is identical for all neurons. Hence
$\mu^{(q)}_{t,l}=\mu\left(\gamma^{(0)}_{t,l}+\gamma^{(1,q)}_{t,l}\right)$,
with neuron-dependent corrections $\gamma^{(1,q)}_{t,l}$ coming from the rank-1 gate sensitivity and candidate path in Eq. 53.
(c) Constant scalar gate $s\in(0,1)$ (“ConstGate”)
As a minimal baseline we fix the gate value to a constant scalar $s\in(0,1)$, independent of $t$, inputs, or hidden state. The update is
$h_{t}\;=\;(1-s)\,h_{t-1}\;+\;s\,\tilde{h}_{t},\qquad\tilde{h}_{t}=\tanh(a^{h}_{t}),\;\;a^{h}_{t}=W_{h}x_{t}+U_{h}h_{t-1}+b_{h}.$ (54)
The Jacobian is
$J_{t}\;=\;\underbrace{(1-s)\,I}_{\text{fixed leak}}\;+\;\underbrace{s\,S^{h}_{t}\,U_{h}}_{\text{candidate path}}.$ (55)
Unlike DiagGate and SharedGate, there are no gate sensitivities because $s$ is constant and not learned. The dynamics are a rigid combination of identity and the candidate path, with a fixed leakage rate (1s)(1-s) applied to all neurons.
Effective learning rates.
Because $s$ is constant, the envelope is trivial:
$\gamma^{(0)}_{t,l}=(1-s)^{t-l},$
identical for all neurons. Thus
$\mu^{(q)}_{t,l}=\mu\left((1-s)^{t-l}+\gamma^{(1,q)}_{t,l}\right)$,
with $\gamma^{(1,q)}_{t,l}$ depending only on the candidate-path correction.
Appendix C Noise floor of the empirical matched statistic
This appendix clarifies the assumptions underlying the noise model used in the manuscript and justifies the scaling of stochastic fluctuations in the empirical matched statistic. We distinguish carefully between exact distributional assumptions and domain-of-attraction arguments, and emphasize the role of linearity in preserving heavy-tailed behavior [35].
Empirical averaging over independent sequences.
By construction, the empirical matched statistic introduced in Section 5.2 at lag $\ell$ is an average over $N$ independently generated training sequences,
$\widehat{T}_{N}(\ell)=\frac{1}{N}\sum_{n=1}^{N}T^{(n)}_{t,\ell}$,
where $T^{(n)}_{t,\ell}$ denotes the matched statistic computed from the $n$th sequence. Independence across sequences implies that the summands $\{T^{(n)}_{t,\ell}\}_{n=1}^{N}$ are independent for fixed $(t,\ell)$.
Heavy-tailed fluctuations and domain of attraction.
Empirical studies of stochastic gradients in deep neural networks consistently report heavy-tailed fluctuations, with tail index $\alpha\in(1,2)$ in many practical settings. Accordingly, we assume that the centered per-sequence matched statistic
$T^{(n)}_{t,\ell}-\mathbb{E}[T_{t,\ell}]$
belongs to the domain of attraction of a symmetric $\alpha$-stable law. This assumption does not require the statistic to be exactly $\mathrm{S}\alpha\mathrm{S}$ distributed at finite sample sizes; it only requires regularly varying tails with index $\alpha$.
Linearity and preservation of heavy tails.
The matched statistic $T_{t,\ell}$ is a finite linear combination of gradient components; see Eq. (29). If the underlying gradient fluctuations have heavy tails with index $\alpha$, then each term $\zeta^{(q)}_{t,\ell}$, being a linear functional of the gradient, belongs to the same domain of attraction. Finite linear combinations preserve the tail index, so $T_{t,\ell}-\mathbb{E}[T_{t,\ell}]$ also lies in the domain of attraction of a symmetric $\alpha$-stable law, with a scale parameter $\sigma_{\alpha}(\ell)$ that may depend on the lag $\ell$.
Canonical $\alpha$-stable representation.
For concreteness and analytical tractability, we represent the centered statistic by a canonical element of its domain of attraction,
$T^{(n)}_{t,\ell}-\mathbb{E}[T_{t,\ell}]\sim\mathrm{S}\alpha\mathrm{S}\!\big(0,\sigma_{\alpha}(\ell)\big).$
This modeling choice serves as a representative of the entire domain of attraction: all results derived below depend only on the tail index $\alpha$ and the scaling behavior of sums and averages, not on the exact finite-sample distribution.
Scaling of the empirical average.
A defining property of symmetric $\alpha$-stable laws is stability under summation. If $\{X_{n}\}_{n=1}^{N}$ are independent $\mathrm{S}\alpha\mathrm{S}(0,\sigma)$ random variables, then
$\sum_{n=1}^{N}X_{n}\sim\mathrm{S}\alpha\mathrm{S}\!\big(0,\sigma N^{1/\alpha}\big).$
Because the centered matched statistics belong to the same domain of attraction, their sum obeys the same scaling, yielding
$\sum_{n=1}^{N}\big(T^{(n)}_{t,\ell}-\mathbb{E}[T_{t,\ell}]\big)\sim\mathrm{S}\alpha\mathrm{S}\!\big(0,\sigma_{\alpha}(\ell)N^{1/\alpha}\big).$
Dividing by $N$ gives the empirical average
$\widehat{T}_{N}(\ell)-\mathbb{E}[T_{t,\ell}]\sim\mathrm{S}\alpha\mathrm{S}\!\big(0,\sigma_{\alpha}(\ell)N^{1/\alpha-1}\big)=\mathrm{S}\alpha\mathrm{S}\!\Big(0,\frac{\sigma_{\alpha}(\ell)}{N^{\,1-1/\alpha}}\Big).$
Noise floor and concentration rate.
The typical magnitude of stochastic fluctuations in the empirical matched statistic therefore decays as $N^{-(1-1/\alpha)}$. For $\alpha=2$ this recovers the classical Gaussian rate $N^{-1/2}$, while for $\alpha<2$ concentration is slower due to heavy-tailed noise. This $N^{-(1-1/\alpha)}$ noise floor underlies the detectability thresholds and sample-complexity bounds derived in the manuscript and determines how heavy-tailed gradient statistics compress the learnability window.
Appendix D LAN-based KL bound for averaged $\alpha$-stable location models
For completeness, we derive the KL lower bound used in Section 5.3.2 for the empirical matched statistic $\widehat{T}_{N}(\ell)$, which is defined as an average over $N$ independent training sequences. Unlike the classical LAN setting for product measures, the present analysis concerns a shrinking-scale $\alpha$-stable location family. The argument follows the general LAN framework for triangular arrays of statistical experiments with vanishing noise levels [45].
Statistical model and parameterization.
We consider the statistical experiment induced by the empirical matched statistic $\widehat{T}_{N}(\ell)$ at a fixed lag $\ell$, which is defined as an average over $N$ independent training sequences. As shown in Appendix C, under heavy-tailed gradient fluctuations with tail index $\alpha>1$, the centered statistic $\widehat{T}_{N}(\ell)-\mathbb{E}[T_{t,\ell}]$ exhibits stochastic fluctuations whose typical magnitude scales as $N^{-(1-1/\alpha)}$.
To formalize hypothesis testing at lag $\ell$, we model the distribution of the centered empirical statistic by a location family with shrinking scale. Specifically, let $Z$ be a symmetric $\alpha$-stable random variable with unit scale and characteristic exponent $\alpha>1$, and define the observation
$Y_{N}=\theta+s_{N}Z,\qquad s_{N}:=\sigma_{\alpha}(\ell)\,N^{1/\alpha-1}.$ (56)
Here, $Y_{N}$ represents the random fluctuation of the empirical matched statistic, and $\theta\in\mathbb{R}$ is a location parameter encoding the presence or absence of a lag-$\ell$ dependency.
In the detection problem considered in the main text, $\theta$ corresponds to the mean shift induced by the effective learning-rate alignment:
$\theta\in\left\{+\tfrac{1}{2}\Delta(\ell),\;-\tfrac{1}{2}\Delta(\ell)\right\},\qquad\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell).$
Thus, the two hypotheses differ only by a small location perturbation whose magnitude is controlled by the effective learning-rate envelope.
For $\alpha>1$, symmetric $\alpha$-stable laws are absolutely continuous and admit a density. Denoting by $P_{\theta,N}$ the law of $Y_{N}$, its density is given by
$p_{\theta,N}(y)=\frac{1}{s_{N}}\,f_{0}\!\left(\frac{y-\theta}{s_{N}}\right),$ (57)
where $f_{0}$ is the density of a fixed symmetric $\alpha$-stable distribution with unit scale.
Triangular-array structure and relation to classical LAN.
Unlike the classical LAN framework, which considers $N$ i.i.d. observations from a fixed distribution, the present setting involves a triangular array of statistical experiments. This means that for each sample size $N$ we observe a single random variable $Y_{N}$, whose distribution depends on $N$ through the shrinking scale $s_{N}$.
Intuitively, this structure arises because $\widehat{T}_{N}(\ell)$ is itself an average over $N$ independent sequences: increasing $N$ does not provide more independent observations of a fixed noise model, but instead produces a more concentrated version of the same statistic. As $N$ grows, the noise scale $s_{N}$ tends to zero, yielding a sequence of location families with vanishing dispersion.
Such shrinking-scale location models are a standard object in asymptotic statistics [45] and admit local asymptotic normality under mild regularity conditions. In the present case, for $\alpha>1$, the family $\{P_{\theta,N}\}_{N\geq 1}$ satisfies a LAN expansion with normalization rate $r_{N}=s_{N}^{-1}$.
LAN for shrinking-scale location families.
For $\alpha>1$, symmetric $\alpha$-stable densities possess finite Fisher information for the location parameter [20, 26, 35, 45]. As $s_{N}\to 0$, the family $\{P_{\theta,N}\}_{N\geq 1}$ forms a triangular array of location models that is locally asymptotically normal with normalization rate
$r_{N}:=\frac{1}{s_{N}}=\frac{N^{\,1-1/\alpha}}{\sigma_{\alpha}(\ell)}.$
Specifically, there exists a constant $I_{\alpha}>0$, depending only on the tail index $\alpha$, and a sequence of random variables $\Delta_{N}$ such that, for each fixed $h\in\mathbb{R}$,
$\log\frac{dP_{\theta+h/r_{N},N}}{dP_{\theta,N}}=h\,\Delta_{N}-\tfrac{1}{2}h^{2}I_{\alpha}+o_{P_{\theta,N}}(1),\qquad\Delta_{N}\xrightarrow{d}\mathcal{N}(0,I_{\alpha}),$ (58)
as $N\to\infty$. The constant $I_{\alpha}$ plays the role of an effective information coefficient in the local asymptotic expansion and does not require an explicit expression.
Taking expectations with respect to $P_{\theta,N}$ and using $\mathbb{E}[\Delta_{N}]=0$ yields the asymptotic expansion of the KL divergence [26]:
$D_{\mathrm{KL}}\!\big(P_{\theta+h/r_{N},N}\|P_{\theta,N}\big)=\tfrac{1}{2}h^{2}I_{\alpha}+o(1).$ (59)
Consequently, for all sufficiently large $N$, there exists a constant $c_{\alpha}>0$ such that
$D_{\mathrm{KL}}\!\big(P_{\theta+h/r_{N},N}\|P_{\theta,N}\big)\;\geq\;c_{\alpha}\,h^{2}.$ (60)
Application to the detection problem in the main text.
In the main text, the detection and non-detection hypotheses correspond to two symmetric location parameters separated by a fixed amount $\Delta(\ell)$:
$\theta_{1}=+\tfrac{1}{2}\Delta(\ell),\qquad\theta_{0}=-\tfrac{1}{2}\Delta(\ell),$
so that $\theta_{1}-\theta_{0}=\Delta(\ell)$.
The empirical matched statistic has a shrinking noise scale $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$ (Appendix C). To express the fixed separation in the LAN framework, we measure it in units of this scale and define the normalized separation
$h_{N}:=\frac{\theta_{1}-\theta_{0}}{s_{N}}=\frac{N^{\,1-1/\alpha}\,\Delta(\ell)}{\sigma_{\alpha}(\ell)}.$
Substituting $h=h_{N}$ into the LAN-based KL lower bound (60) yields
$D_{\mathrm{KL}}\!\left(P_{\text{det}}\|P_{\text{non}}\right)\;\geq\;c_{\alpha}\,h_{N}^{2}=c_{\alpha}\,\frac{N^{\,2(1-1/\alpha)}\,\Delta(\ell)^{2}}{\sigma_{\alpha}(\ell)^{2}},$ (61)
which is exactly the bound used in Eq. (33) of the main text.
Appendix E From KL divergence to mutual information and Fano bounds
This appendix provides a self-contained derivation of the mutual-information lower bound and the Fano-type sample complexity inequality used in Sec. 5.3.2. The arguments rely only on standard relations between KL divergence, mutual information, and binary hypothesis testing [11].
Binary detection setup.
Fix a lag $\ell$. The hypotheses in Sec. 5.3.2 correspond to two $\alpha$–stable location models for the empirical matched statistic $T=\widehat{T}_{N}(\ell)$, obtained by averaging over $N$ independent training sequences. Introduce a binary label
$B\in\{0,1\},\qquad\mathbb{P}(B{=}1)=\mathbb{P}(B{=}0)=\tfrac{1}{2},$
where B=1B{=}1 selects the detection distribution $P_{\text{det}}$ and B=0B{=}0 selects the non-detection distribution $P_{\text{non}}$. The mutual information between $B$ and $T$ is defined as
$I(B;T)=D_{\mathrm{KL}}\!\big(P_{B,T}\,\|\,P_{B}P_{T}\big),$
where $P_{B,T}$ denotes the joint distribution and $P_{B}P_{T}$ the product of the marginals.
Mutual information via the mixture identity.
Because $B$ is equiprobable, the joint density factorizes as
$p_{B,T}(b,t)=p_{B}(b)\,p_{T\mid B}(t\mid b)=\tfrac{1}{2}\,p_{T\mid B}(t\mid b),$
with
$p_{T\mid B}(t\mid 1)=p_{\text{det}}(t),\qquad p_{T\mid B}(t\mid 0)=p_{\text{non}}(t).$
The marginal distribution of $T$ is the mixture
$M=\tfrac{1}{2}P_{\text{det}}+\tfrac{1}{2}P_{\text{non}}.$
Starting from the definition of mutual information,
$I(B;T)=\sum_{b=0}^{1}\int p_{B}(b)\,p_{T\mid B}(t\mid b)\log\frac{p_{T\mid B}(t\mid b)}{p_{T}(t)}\,dt,$
one obtains the mixture identity
$I(B;T)=\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{det}}\|M\big)+\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{non}}\|M\big).$ (62)
Since KL divergences are nonnegative,
$I(B;T)\;\geq\;\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{det}}\|M\big),$
and since mutual information cannot exceed the entropy of the binary label, $I(B;T)\leq\log 2$. Combining these observations yields the general lower bound
$I(B;T)\;\geq\;\min\!\left\{\log 2,\;\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{det}}\|M\big)\right\}.$ (63)
In the main text, the factor $\tfrac{1}{2}$ and the difference between $D_{\mathrm{KL}}(P_{\text{det}}\|M)$ and $D_{\mathrm{KL}}(P_{\text{det}}\|P_{\text{non}})$ are absorbed into the constant $c_{\alpha}$, since only the scaling with respect to $N$ and $\Delta(\ell)/\sigma_{\alpha}(\ell)$ is relevant.
Fano’s inequality and sample complexity.
Let $P_{e}$ denote the probability of incorrectly deciding whether a lag-$\ell$ signal is present based on the observation of $T$. For binary hypotheses with equal priors, Fano’s inequality [11] gives
$P_{e}\;\geq\;1-\frac{I(B;T)}{\log 2}.$
Thus, to ensure a target error probability $P_{e}\leq\epsilon<1/2$, it is sufficient that
$I(B;T)\;\geq\;\log 2\,(1-\epsilon).$
Substituting the mutual-information lower bound from Eq. (34) and solving for NN yields the sample complexity inequality
$N\;\geq\;\left(\frac{\sigma_{\alpha}(\ell)}{\sqrt{c_{\alpha}}\,\overline{m}_{\mu}(\ell)\,f(\ell)}\right)^{\!\frac{\alpha}{\alpha-1}}\!\!\Big(\log\tfrac{1}{2\epsilon}\Big)^{\frac{\alpha}{2(\alpha-1)}},$ (64)
which is the bound reported in Sec. 5.3.2.
Appendix F Proofs of Lemmas
Proof of Lemma 5.1
Proof.
Recall that for the diagonally gated models in Appendix B we write $\mu^{(q)}_{t,\ell}=\mu\big(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\big)$, while for the GRU we additionally track the diagonal zeroth-order envelopes $\gamma^{(0,q)}_{t,\ell},\rho^{(0,q)}_{t,\ell},\eta^{(0,q)}_{t,\ell}$ (see Eq. (26)); the LSTM zeroth-order coefficient is given by Eq. (13). In all cases, the zeroth-order diagonal envelopes are products of time-local gate factors and bounded activation derivatives, hence lie in $[0,1]$.
Since the first-order diagonal correction may contain signed diagonal contributions (e.g. through diagonal entries of recurrent matrices), we prove monotonicity for the absolute diagonal envelope used throughout the empirical pipeline:
[…]
Zeroth order.
For the diagonally gated RNNs in Appendix B, for each neuron $q$ the zeroth-order term has the form
$\gamma^{(0,q)}_{t,\ell}\;=\;\prod_{j=t-\ell+1}^{t}a^{(q)}_{j},\qquad a^{(q)}_{j}\in[0,1],$
so that $\gamma^{(0,q)}_{t,\ell+1}=\gamma^{(0,q)}_{t,\ell}\,a^{(q)}_{t-\ell}\leq\gamma^{(0,q)}_{t,\ell}$. For the GRU, the additional envelopes $\rho^{(0,q)}_{t,\ell}=\prod_{j=t-\ell+1}^{t}r_{j,q}$ and $\eta^{(0,q)}_{t,\ell}=\prod_{j=t-\ell+1}^{t}(1-z_{j,q})\,r_{j,q}$ obey the same multiplicative extension property and are therefore nonincreasing in $\ell$. For the LSTM, Eq. (13) gives the single surviving zeroth-order path
$\gamma^{(0,q)}_{t,\ell}=e_{t,q}\prod_{j=\ell+1}^{t}f_{j,q},\qquad e_{t,q}\in[0,1],\;f_{j,q}\in[0,1],$
which is again nonincreasing as $\ell$ increases.
First-order diagonal correction.
Fix $q$. The diagonal first-order term is a finite sum of contributions indexed by the insertion time $p$ (and, in general, by a finite set of diagonal factors produced by the architecture at time $p$). Each such contribution has the form
$C^{(q)}_{p,t,\ell}\;=\;b^{(q)}_{p}\,\prod_{j\in\mathcal{I}_{p,\ell}}a^{(q)}_{j},\qquad a^{(q)}_{j}\in[0,1],$
where $b^{(q)}_{p}$ is $\ell$-independent (possibly signed), and where the index sets satisfy $\mathcal{I}_{p,\ell}\subseteq\mathcal{I}_{p,\ell+1}$ (i.e. increasing the lag can only add further multiplicative gate factors). Therefore,
[…]
Conclusion.
By the triangle inequality,
[…]
Proof of Lemma 5.2
Proof.
Starting from the per-lag sufficient sample size in Eq. (39),
$N(\ell)=\kappa_{\alpha,\epsilon}\,\left(\frac{\sigma_{\alpha}(\ell)}{\overline{m}_{\mu}(\ell)\,f(\ell)}\right)^{\!\kappa_{\alpha}},\qquad\kappa_{\alpha,\epsilon}=\frac{1}{c_{\alpha}^{\kappa_{\alpha}/2}}\Big(\log\tfrac{1}{2\epsilon}\Big)^{\kappa_{\alpha}/2},$
and applying the boundedness assumptions $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$ and $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$, we obtain
$\kappa_{\alpha,\epsilon}\left(\frac{c_{\sigma}}{C_{m}}\right)^{\!\kappa_{\alpha}}\,f(\ell)^{-\kappa_{\alpha}}\ \leq\ N(\ell)\ \leq\ \kappa_{\alpha,\epsilon}\left(\frac{C_{\sigma}}{c_{m}}\right)^{\!\kappa_{\alpha}}\,f(\ell)^{-\kappa_{\alpha}},$
which yields Eq. (40) with $c_{\star}=\kappa_{\alpha,\epsilon}(c_{\sigma}/C_{m})^{\kappa_{\alpha}}$ and $C_{\star}=\kappa_{\alpha,\epsilon}(C_{\sigma}/c_{m})^{\kappa_{\alpha}}$.
For the horizon bound, recall the detectability condition from Eq. (37):
$f(\ell)\geq\ \frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)}.$
As before, the constant factor $\frac{1}{\sqrt{c_{\alpha}}}\sqrt{\log\tfrac{1}{2\epsilon}}$ is independent of $\ell$ and $N$ and can be absorbed into the constants of the sandwich bound.
Bounding the right-hand side above and below using the assumed bounds on $\sigma_{\alpha}(\ell)$ and $\overline{m}_{\mu}(\ell)$ gives
$\frac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\ \leq\ \frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)}\ \leq\ \frac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}.$
The detectability condition is guaranteed whenever
$f(\ell)\geq\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}},$
and cannot hold unless
$f(\ell)\geq\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}.$
Since $f(\ell)$ is nonincreasing (Lemma 5.1), the admissible lags form an interval $0,f^{\leftarrow}(x)]$. Applying the generalized inverse $f^{\leftarrow}$ yields
$f^{\leftarrow}\!\Big(\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}\Big)\ \leq\ \mathcal{H}_{N}\ \leq\ f^{\leftarrow}\!\Big(\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\Big),$
which establishes the claimed sandwich bound. ∎
Remark.
The generalized inverse is used because $f(\ell)$ need not be strictly decreasing. By Lemma 5.1, $f(\ell)$ is nonincreasing, ensuring that $f^{\leftarrow}$ is well defined. The proof shows that the dependence on stochasticity and alignment enters only through the constants $(c_{\sigma},C_{\sigma},c_{m},C_{m})$ and the tail index $\alpha$ (via $\kappa_{\alpha}$); the functional dependence of $N(\ell)$ and $\mathcal{H}_{N}$ on the envelope $f(\ell)$ is dictated entirely by the power $f(\ell)^{-\kappa_{\alpha}}$.
On the boundedness assumptions.
The conditions $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$ and $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$ should be understood as mild regularity assumptions over the finite range of lags under consideration. Absent pathological degeneracies, such as vanishing alignment $\overline{m}_{\mu}(\ell)\to 0$ or diverging fluctuation scale $\sigma_{\alpha}(\ell)\to\infty$, these quantities remain finite and vary smoothly with $\ell$. Thus, the boundedness assumptions affect only constant factors in Eq. (40), while the asymptotic dependence of $N(\ell)$ and $\mathcal{H}_{N}$ on the envelope $f(\ell)$ is entirely captured by the exponent $\kappa_{\alpha}$.
Appendix G Asymptotic Scaling of the Learnability Window under $\alpha$-Stable Noise
This appendix provides the detailed derivations supporting the scaling relations summarized in Sec. 5.4 and established by Lemmas 5.15.2. Throughout, we use the corrected concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$, which governs the statistical behavior of empirical averages under $\alpha$-stable noise. Starting from the $\alpha$-stable model of Eq. (32) and the finite-sample bound of Eq. (35), we show how the slow $N^{1/\kappa_{\alpha}}$ concentration rate characteristic of heavy-tailed averages determines the asymptotic growth of the learnability window $\mathcal{H}_{N}$.
Per-lag sufficient sample size.
From Eq. (35), the minimal number of independent sequences required to detect a lag-$\ell$ dependency with error probability $P_{e}\!\leq\!\epsilon$ is
$N(\ell)\;=\;\kappa_{\alpha,\epsilon}\,\left(\frac{\sigma_{\alpha}(\ell)}{\overline{m}_{\mu}(\ell)f(\ell)}\right)^{\!\kappa_{\alpha}},\qquad\kappa_{\alpha,\epsilon}=\frac{1}{c_{\alpha}^{\kappa_{\alpha}/2}}\,\Big(\log\tfrac{1}{2\epsilon}\Big)^{\kappa_{\alpha}/2}.$ (65)
The constant $\kappa_{\alpha,\epsilon}$ depends only on the tail index $\alpha$ and the target detection error $\epsilon$, and can be absorbed into the asymptotic constants below. We assume the boundedness conditions of Lemma 5.2,
$c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma},\qquad c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m},$
and recall that $\ell\mapsto f(\ell)$ is nonincreasing (Lemma 5.1).
Two-sided sandwich bound.
Substituting these bounds into (65) yields constants
$c_{\star}:=\kappa_{\alpha,\epsilon}\Big(\tfrac{c_{\sigma}}{C_{m}}\Big)^{\!\kappa_{\alpha}},\qquad C_{\star}:=\kappa_{\alpha,\epsilon}\Big(\tfrac{C_{\sigma}}{c_{m}}\Big)^{\!\kappa_{\alpha}},$
such that
$c_{\star}\,f(\ell)^{-\kappa_{\alpha}}\ \leq\ N(\ell)\ \leq\ C_{\star}\,f(\ell)^{-\kappa_{\alpha}}.$ (66)
Hence, $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, formalizing that longer dependencies (smaller $f(\ell)$) require increasingly large sample sizes to detect under heavy-tailed noise.
Learnability window as a level set.
From the finite-sample detectability condition of Eq. (37),
$f(\ell)\ \geq\ \frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)},$
and using the same bounding arguments, we can express the learnability window $\mathcal{H}_{N}$ as a level set of $f(\ell)$:
$f^{\leftarrow}\!\Big(\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}\Big)\ \leq\ \mathcal{H}_{N}\ \leq\ f^{\leftarrow}\!\Big(\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\Big),$ (67)
where $f^{\leftarrow}(y)=\sup\{\ell\geq 1:f(\ell)\geq y\}$ denotes the generalized inverse. By Lemma 5.1, $f(\ell)$ is nonincreasing, ensuring that the generalized inverse $f^{\leftarrow}$ is well-defined. This reproduces the sandwich bound in Eq. (41) of the main text.
Asymptotic regimes.
We now invert (66) (equivalently (67)) for three canonical decay laws of the envelope $f(\ell)$.
(i) Logarithmic decay.
If $f(\ell)\asymp c/\log(1+\ell)$ with $c>0$, then
$N(\ell)\asymp[\log(1+\ell)]^{\kappa_{\alpha}}\quad\Longrightarrow\quad\log(1+\ell)\asymp N^{1/\kappa_{\alpha}}\quad\Longrightarrow\quad\mathcal{H}_{N}\asymp\exp(\kappa N^{1/\kappa_{\alpha}})-1.$
(ii) Polynomial (algebraic) decay.
If $f(\ell)\asymp c\,\ell^{-\beta}$ with $\beta>0$, then
$N(\ell)\asymp\ell^{\kappa_{\alpha}\beta}\quad\Longrightarrow\quad\mathcal{H}_{N}\asymp N^{1/(\kappa_{\alpha}\beta)}.$
(iii) Exponential (geometric) decay.
If $f(\ell)\asymp c\,\lambda^{\ell}$ with $\lambda\in(0,1)$, then
$N(\ell)\asymp\lambda^{-\kappa_{\alpha}\ell}\quad\Longrightarrow\quad\mathcal{H}_{N}\asymp\frac{\log N}{\kappa_{\alpha}\,\log(1/\lambda)}.$
All asymptotic forms hold up to multiplicative constants inherited from $c_{\sigma}$, $C_{\sigma}$, $c_{m}$, $C_{m}$, and $\kappa_{\alpha,\epsilon}$. As $\alpha$ decreases, heavier tails increase $\kappa_{\alpha}$ and slow the $N^{1/\kappa_{\alpha}}$ concentration of empirical averages, uniformly compressing the learnability window $\mathcal{H}_{N}$ across all decay regimes.